OpenStudy (anonymous):

A motorboat travels 384km in 6 hours going upstream and 408 km in 4 hours going downstream. What is the rate of the boat in still water and what is the rate of the current? PLEASE HELP

OpenStudy (anonymous):

Please Help.

OpenStudy (anonymous):

Let x be the rate of the boat in still water and y be the rate of the current. The rates are distance over time For upstream subtract the current. x - y = 384/6 For downstream add the current x + y = 408/4

OpenStudy (anonymous):

I understand the concept but I'm still getting the wrong answers, could you show me what you got and how you go it please?

OpenStudy (anonymous):

x - y = 64 x + y = 102 Add the equations together to eliminate y 2x = 166 x = 83 km/h Then substitute to get y.